OpenStudy (anonymous):

A sample of Freon-12 (CF2Cl2) occupies 25.5 L at 298 K and 153.3 kPa. Find its volume at STP.

OpenStudy (jfraser):

do you know the combined gas law?

OpenStudy (anonymous):

i know all of that stuff, PV=nRT but what is STP, (my textbooks always throwing these curve balls right when i think i get it lol)

OpenStudy (jfraser):

since gases can be used at all sorts of temperatures and pressures, we use STP to mean: standard temperature and pressure: 273K and 1.00atm. That way we can compare different gases under identical conditions

OpenStudy (anonymous):

ok so 273 is the standard temp and 1.00 atm is the standard pressure, how do i compare the 1 atm to the 153.3 kpa? I would want to multiply by 760 but thats just the only idea i have and it seems really wrong.

OpenStudy (jfraser):

760 is the conversion for mmHg, not kPa. the standard measurements of pressure are:

OpenStudy (jfraser):

\[1atm = 760mmHg \space (or \space Torr) = 101.3kPa = 14.7PSI\]

OpenStudy (jfraser):

so divide 153kPa by 101.3, like so:\[153.3\cancel{kPa} * \frac{1atm}{101.3\cancel{kPa}}\]

OpenStudy (jfraser):

see how the units disappear and we just do the math of the numbers that are left behind? As a general rule, you always want to look at the units first, and the numbers have to come along for the ride, so long as the units are put in the right places first.

OpenStudy (anonymous):

yeah thats usually my mentality when i'm doing these problems. "Ok i need volume, if my answer ends up having the units for volume it's probably right". Now that the pressure is all cleared up i have a couple more questions, one being what PSI stood for and should i stress knowing that?(which is really an irrelevant question i'm just curious). My real question is will my answer come from comparing the ratios from the "standard" to the given in order to find the volume.

OpenStudy (jfraser):

PSI is pounds per square inch, which is an americanized pressure instead of kPa

OpenStudy (jfraser):

your answer will come from plugging the values into the combined gas law, since you have 5 of the 6 pieces. all that is missing is the second volume\[\frac{P_1*V_1}{T_1} = \frac{P_2*V_2}{T_2}\]

OpenStudy (anonymous):

we don't care about the n1 and n2 ( i only say that because it's on the formula sheet i have in front of me, where right next to T1 and T2 are N1 and N2)

OpenStudy (jfraser):

since we're not told how many moles are in the container, the value for \(n_1\) and \(n_2\) must be the same, whatever they are. So they cross out of the equation

OpenStudy (anonymous):

gotcha man, one last thing (gonna milk this free tutoring session, but no worries i got you on the "awesome tutor" rating thing in the corner =p ) we have 2 formulas , the fractional one that you showed above including n1 and n2 and then you have Pv=nRT, is there certain scenarios where one formula is more appropriate than the other, or using one formula may seem like the right way but gives you a totally different(and wrong) answer.

OpenStudy (anonymous):

I saw you typing and stopped lol, it's cool i appreciate all the help JFraser and you deserve them best response points, thank you!

OpenStudy (jfraser):

if you're given one set of conditions, and then a DIFFERENT set of conditions, then you're looking to use the combined law. If you're given a P, V, n, and\or T and asked to find the missing one, then you're using the ideal-gas law

OpenStudy (anonymous):

If i could hit best response twice i would, thank you so much for your help.

OpenStudy (jfraser):

yw